Solve using substitution process.
15x-9y=-3
-5x+3y=-7

Answer by richard1234(7193)

No solutions. You can multiply the second equation by 3 to obtain

15x - 9y = -3
-15x + 9y = -21

We see that if there is a solution (x,y), then 0 = -24 if we add the two equations.